OpenStudy (anonymous):

What is the slope of a line that passes through the point (−5, 3) and is parallel to a line that passes through (2, 13) and (−4, −11)?

OpenStudy (nincompoop):

first, identify the slope of the second line with (2,13) (-4, -11)

OpenStudy (anonymous):

I forgot how to do that...

OpenStudy (nincompoop):

\[slope = \frac{\Delta y}{\Delta x}=\frac{y_{final}-y_{initial}}{x_{final}-x_{initial}}\]

OpenStudy (nincompoop):

do you recall now?

OpenStudy (anonymous):

Yes

OpenStudy (nincompoop):

okay solve for the slope of the second line

OpenStudy (only4u):

slope of two parallel lines are same

OpenStudy (anonymous):

So I do 13-11/2-4?

OpenStudy (nincompoop):

@only4u you can move on to a different question now, I got this

OpenStudy (nincompoop):

I think you're forgetting the proper signs, candee

OpenStudy (only4u):

it will be 4

OpenStudy (nincompoop):

-11 and -4 correct? \[slope = \frac{(13)- (-11)}{(2) - (-4)}\]

OpenStudy (nincompoop):

see how I didn't leave out the signs ?

OpenStudy (anonymous):

Okay so after I subtract -11 and 4 will I have my answer?

OpenStudy (nincompoop):

you will get your slope

OpenStudy (nincompoop):

@only4u is itching to teach you properly

OpenStudy (nincompoop):

he is frustrated that you are not learning fast enough

OpenStudy (only4u):

yes :) exactly :)

OpenStudy (anonymous):

So the answer is 4

OpenStudy (nincompoop):

okay you can leave now, we want to do this slowly and precisely

OpenStudy (nincompoop):

okay slope of the second line is 4 NOTE that parallel lines have the same slope so this implies that the first line with the point (-5, 3) has the same slope do you accept this?

OpenStudy (anonymous):

Yes! Thank you so much for your help

OpenStudy (only4u):

u r always welcome candee :)

OpenStudy (nincompoop):

okay so there are two things involved in this concept the relationship of the slope of parallel lines and the relationship of the slope of perpendicular lines so you can move on to your other slope question BEAR in mind that perpendicular lines have negative reciprocal slopes what do I mean? when one of the lines have a slope of 3 (the same as 3/1), the other line that is perpendicular would have a slope of -1/3 another example when one of the lines have a slope of 1/5, then the other line that is perpendicular would have a slope of -5 REMEMBER: negative reciprocal

OpenStudy (nincompoop):

are you taking notes?

OpenStudy (triciaal):

@only4u "Don't just provide the answer to a problem when someone else is in the middle of helping! But if you want to help, by all means, join in!"